# UniGLM: Training One Unified Language Model for Text-Attributed Graphs

### v1

@@ 4 Experiments
Throughout the experiments, we aim to answer the following research questions. RQ1: How does UniGLM perform against leading graph embedding models in terms of node classification and link prediction tasks? RQ2: How well does UniGFM transfer in cross-domain and in-domain scenarios? RQ3: How does each component of UniGLM, i.e., sampling strategy and efficient embedding table, contribute to the performance? RQ4: What is the impact of different pre-trained language model backbones and hyper-parameter on UniGLM?
4.1 Experiments Setup
@@
We evaluate UniGLM on eight TAG datasets (details in Tabel  9 in Appendix). UniGLM is compared with multiple embedding models for node classification and link prediction. Node classification is conducted with MLP and GNNs (GCN, GraphSAGE, RevGAT), while link prediction uses MLP and Graph AutoEncoders. Our experiments focus on semi-supervised and transfer learning settings. More details can be found in Appendix.
4.2 Performance Evaluation
Node Classification. To answer RQ1, we conduct extensive experiments on eight benchmark TAG datasets for node classification under the semi-supervised setting. The results are shown in Table 1 and in Table 8(Appendix). We observe that: ① UniGLM significantly outperforms existing graph embedding models in node classification task with state-of-the-art GNNs and MLP. Table 1 demonstrates that UniGLM consistently achieves superior performance across most datasets and models, often ranking first. Although GIANT and PATTON improve over the SE method by leveraging language models and graph structure, UniGLM excels in most cases. ② Performance improvements are proportional to the volume of data used in training, demonstrating the effectiveness of the scaling law for TAGs. Specifically, as depicted in Figure 2, training UniGLM across all TAGs consistently outperforms the variant OneGLM trained on only a single TAG, with the largest observed performance gap being 10%. Furthermore, as indicated in Table 1 and Table 8, co-purchase networks exhibit more substantial improvements compared to citation networks.
Link Prediction. To address RQ1, we evaluate the UniGLM model for link prediction tasks using a 5% training edge set under a transfer learning setting. We train UniGLM on citation networks and evaluate on co-purchase networks. The results are illustrated in Figure 3 and Figure 9 (Appendix). We observe that ③UniGLM exhibits a robust cross-domain transfer capability, generally surpassing other baselines in link prediction. Specifically, UniGLM achieves over $90\%$ on Ogbn-Product dataset, surpassing SE and BERT over $15\%$.
Figure 2: Comparison of UniGLM and OneGLM with node classification task on GCN.
Figure 3: Comparison of UniGLM and baselines with Link prediction task on MLP.
4.3 Transfer Ability
We discuss RQ2 in this section. From the cross-domain perspective, we train model on one domain and apply the model to another. For instance, we train on co-purchase datasets and test on citation networks, or vice versa. These setups are benchmarked against BERT, GIANT fine-tuned on Computers, and Patton pre-trained on Ogbn-Arxiv and evaluate with node classification task. ④We observe that training UniGLM with one domain can enhance the performance of other domain. Specifically, in Table 2 on History, UniGLM trained on citation networks improves performance most in co-purchase networks comparing to baselines, achieving a notable 80.22% accuracy.
Dataset Emb Type Accuracy
History BERT 78.79 ± 0.31
GIA(Computers) 77.35 ± 0.89
PATTON(Arxiv) 79.78 ± 0.29
UniGLM(Citation) 80.22 ± 0.40
Photo BERT 60.03 ± 0.14
GIA(Computers) 75.37 ± 0.48
PATTON(Arxiv) 62.53 ± 0.68
UniGLM(Citation) 63.81 ± 0.36
PubMed BERT 59.79 ± 2.71
GIA(Computers) 57.88 ± 4.56
PATTON(Arxiv) 61.37 ± 1.71
UniGLM(Purchase) 68.29 ± 3.63
Table 2: Accuracy under MLP with Cross-Domain Embeddings. Embedding types under transfer learning settings(cross-domain) are underlined for comparison. Highest results are bolded. Notebly, we do not consider GIA(Computers) for Photo as cross-domain senario for they are both co-purchase network.
For RQ2 in-domain perspective, we employ UniGLM, trained using eight TAGs as the text encoder, to encode the unseen VideoGames dataset, evaluating on both tasks. We use co-viewed edges for training and evaluate on co-purchased edges for link prediction. Results are shown in Table 3. ⑥ We observe that UniGLM exhibits superior in-domain transfer ability. Specifically, UniGLM significantly outperforms BERT on both tasks, highlighting UniGLM’s robustness and effectiveness in in-domain transfer ability.
UniGLM BERT
Link Prediction AP $79.36\pm 0.01$ $58.23\pm 1.27$
AUC $78.39\pm 0.01$ $57.54\pm 1.07$
Node Classification ACC $50.88\pm 0.47$ $50.21\pm 0.28$
@@ 4.3 Transfer Ability
Table 3: In-domain Transfer Ability-Comparison of UniGLM and BERT on unseen VideoGames dataset
4.4 Ablation Study
To answer RQ3, from the perspective of sampling strategies, we employ various methods for selecting positive samples with results shown in Figure 4. ⑦ We observe that it is crucial to consider both node degree and graph density for proper structure information. To further answer RQ3, exploring the effect of embedding table, we compare performance on node classification with or without embedding table in Figure 5 and Figure 8. ⑧ We observe that the embedding table not only decreases the training time but also increases the performance of UniGLM.
Figure 4: Ablation Study - Effect of Sample Strategy.
Figure 5: Ablation Study - Effect of Embedding Table.
To demonstrate the effect of different types of pre-trained LM as backbones in RQ4, we use BERT, RoBERTa and DeBERTa respectively and the results can be seen in Figure 6 and Figure 10. ⑨ We observe that different encoder backbones does not significantly affect the overall performance of UniGLM. The preference for encoders may vary across different datasets. To test the effect of hyper-parameter: number of positive samples in RQ4, we experiment with various sample sizes, and the results are presented in Figure 7, revealing that ⑩ with different sample sizes, results are consistent, UniGLM is not sensitive to this hyper-parameter .
Figure 6: Ablation Study - Effect of different LLM Backbones with node classification task on MLP.
Figure 7: Hyper-parameter Analyse - Number of Positive Samples.
5 Related Work
Our work is related to the following two directions.
Representation Learning on TAGs. Text-attributed graphs (TAGs) have gained significant attention in both academia and industry. Early methods used shallow embedding techniques, which struggled to integrate textual content with graph structure. With pre-trained language models (PLMs), features are now extracted and fed into graph neural networks (GNNs), but this approach often falls short. Recent studies have developed better methods to integrate PLM features into GNNs, improving model performance Ioannidis et al. (2022); Chien et al. (2022); Zhao et al. (2023a); Jin et al. (2023).
Graph Foundation Models. Graph Foundation Models (GFMs) aim to generalize across various graphs and tasks. Xia et al. Xia et al. (2023) proposed OpenGraph, excelling in zero-shot learning. Self-supervised learning (SSL) is crucial for pre-training GFMs. Zhao et al. Zhao et al. (2023b) categorized SSL tasks based on graph-embedded knowledge, enhancing model adaptability. Liu et al. Liu et al. (2023) identified challenges and future directions for GFMs, while Tan et al. Tan et al. (2023) improved GFM generalizability through structure reconstruction with their model.
6 Conclusion
@@
In this study, we introduce UniGLM, a unified framework designed to pre-train a single unified language model for TAGs across domains. UniGLM features two main innovations: (1) an adaptive sampling strategy that selects positive samples and (2) a dynamic embedding table that efficiently encodes these samples on-the-fly to speed up training. We validate UniGLM across diverse TAGs from different domains, where it consistently surpasses existing methods in node classification, link prediction, demonstrating its superior transfer ability, effectiveness and efficiency.
Limitation
In this work, we primarily concentrate on employing language models as unified embedding frameworks for Textual Attributed Graphs (TAGs). Looking ahead, several interesting avenues emerge for extending this research. Firstly, there is a compelling need to explore foundation embedding models tailored for multimodal graphs. These models could integrate diverse types of data, such as textual, visual, and auditory information, enhancing the richness of the embeddings and opening up new possibilities for graph analytics. Secondly, the application of generative language models in graph tasks presents a promising frontier. It is still unknown how our UniGLM can be applied in that direction. These directions not only promise to expand the capabilities of graph neural networks but also bridge the gap between structured graph data and unstructured multimodal data.
Appendix A Appendix
Datasets.  We evaluate the proposed UniGLM framework using eight publicly available TAG datasets. These datasets include two citation networks, namely PubMed Sen et al. (2008) and Ogbn-Arxiv Hu et al. (2020), one co-purchase network Ogbn-Product(subset) from TAPE He et al. (2024) and five E-commerce datasets extracted from Amazon Ni et al. (2019): Electronics-Computers (Computers), Books-History (History), Books-Children (Children), Sports-Fitness (Fitness), and Electronics-Photography (Photo). Noteblely, we use another co-purchase network from He and McAuley (2016) and McAuley et al. (2015) named VideoGames. For node classification, we adhere to the standard data splits used in prior research for PubMed, Ogbn-Arxiv, and Ogbn-Products, while we use a 5:20:75 split for the E-commerce datasets. For link prediction, we adopt the widely used 5:5:90 split and sample equal number of negative links as positive links.
  • PubMed Sen et al. (2008). The PubMed dataset consists of 19,717 scientific publications from PubMed database. The citation network consists of 44,338links
  • Ogbn-Arxiv Hu et al. (2020). The Ogbn-Arxiv dataset is a directed graph, representing the citation network between all Computer Science(CS) arXiv papers.
  • Ogbn-Products(subset) He et al. (2024). The Ogbn-Products dataset represents an Amazon product co-purchasing network, with product decriptions as raw text.
  • Electronics-Computers(Computers) Ni et al. (2019). The Electronics-Computers dataset is a segment of the Amazon co-purchase graph, where nodes represent goods, edges indicate that two goods are frequently bought together, node features are bag-of-words encoded product reviews by default, and class labels are given by the product category.
  • Books-History(History) Ni et al. (2019). Each node represents a book related to history domain, edges indicate that two books are frequently bought together and class labels are given by the book category.
  • Books-Children(Children) Ni et al. (2019). Each node represents a book related to child domain, edges indicate that two books are frequently bought together and class labels are given by the book category.
  • Sports-Fitness(Fitness) Ni et al. (2019). Each node represents a kind of item in Sports and Fitness category.
  • Electronics-Photography(Photo) Ni et al. (2019). This dataset contains review data from Amazon platform. Each node represents a review related to the item in camera category.
Baselines. We compare UniGLM with five textual feature extraction methods. Previously we mentioned four methods including: SE, BERT, GIANT, PATTON in Section 1. To comprehensively compare the proposed UniGLM with existing methods on multiple-datasets, we implement MixGIA as an additional baseline. MixGIA utilizes the GIANT framework and is applied to multiple datasets, leveraging the structural information of each dataset to fine-tune an a shared BERT model. Specifically, original GIANT construct a clustering tree to classify each node layer by layer. Our implementation of MixGIA set a fix number of layers for clustering tree and change the tree alone with the dataset per layer.
Experimental Settings. For node classification, we input the node embeddings from UniGLM into MLP and several different state-of-art GNN models, including GCN Mikolov et al. (2013), GraphSAGE Hamilton et al. (2017) and RevGAT Li et al. (2021). We run experiments 5 times and report the mean result and the standard deviation. We explore link prediction ability under the transfer learning setting, details are explained in link prediction evaluation. We use AUC and AP as metrics for link prediction. We test with MLP and Graph AutoEncoders Kipf and Welling (2016b). For the reproducibility of our experiments, we employ GNN implementations from the PyG (Fey and Lenssen, 2019) package. The default language model backbone is Bert if not specified. Hyper-parameters are shown below.
Hyperparameter Our hyperparameter for GNN is set as same as TAPE He et al. (2024). For UniGLM, we have hyperparameter in Table 4 as follow:
Hyperparameters
BATCH SIZE 64
TEMPERATURE 0.3
MAX SEQUENCE LENGTH 512
NUM POS SAMPLES (t) 6
Table 4: Hyperparemeters of UniGLM
Device and time cost. We compare the training and inference time cost of GIANT, PATTON and UGEmb. For GIANT and PATTON, the training time is the sum of training all 8 TAGs listed above. For UGEmb, it is the time to use all 8 TAGs to train one model. Inference time is the time to encode Ogbn-Arxiv for each method. We use a A800(80G) for training and inference.
GIANT PATTON UGEmb
Training Time 14.24h 5.4d 13.5h
Inference Time 30min 5h15min 29min
Table 5: Time Cost of Different Frameworks
Additional Results In this section, we present further results and details not previously included due to space constraints. These additional findings further substantiate our research claims and provide a deeper understanding of our study’s implications.
This figure  8 shows the result with or without Embedding Table design.
Figure 8: Ablation Study - Effect of Embedding Table.
This figure 9 shows the result of using GAE as the inference model for link prediction.
Figure 9: Average Precision with GAE
This figure 10 shows the result of using different LLMs as backbone.
Figure 10: Ablation Study - Effect of different LLM Backbones with node classification task on GCN.
This table  7 shown additional results to prove the effectiveness of our adaptive sampling strategy.
Dataset Emb Type Accuracy
History with E 82.51±0.08
w/o E 81.96 ± 0.29
Computers with E 81.14±0.19
w/o E 80.14 ± 0.14
Children with E 50.26±0.15
w/o E 49.07 ± 0.25
Fitness with E 90.84±0.08
w/o E 90.68 ± 0.10
Table 6: More results of Ablation Study - Effect of Embedding Table(GCN).
This table  6 shown additional results to prove the effectiveness of our Embedding Table Module.
Dataset Emb Type Accuracy
History w/o Node Degree 79.84 ± 0.24
w/o Graph Density 83.07 ± 0.31
with both 83.42 ± 0.18
PubMed w/o Node Degree 76.04 ± 1.59
w/o Graph Density 82.35 ± 0.83
with both 81.98 ± 1.32
Children w/o Node Degree 47.57 ± 0.29
w/o Graph Density 51.38 ± 0.20
with both 51.86 ± 0.31
Fitness w/o Node Degree 90.03 ± 0.11
w/o Graph Density 88.62 ± 0.07
with both 90.39 ± 0.08
Ogbn-Products w/o Node Degree 74.96 ± 0.57
w/o Graph Density 76.56 ± 1.14
with both 76.46 ± 0.21
Table 7: More results of Ablation Study - Effect of Sample Strategy(MLP).
This table 8 is the other half of node classification results.
Dataset Emb Types MLP GCN SAGE RevGAT
Children SE 38.84±0.35 (+33.52%) 43.19±0.31 (+16.37%) 44.83±0.24 (+16.28%) 43.00±0.18 (+19.40%)
BERT 44.00±0.83 (+17.86%) 46.88±0.60 (+7.21%) 47.97±0.55 (+8.67%) 48.00±0.11 (+6.96%)
GIANT 48.95±0.23 (+5.94%) 48.47±0.35 (+3.69%) 51.41±0.42 (+1.40%) 50.63±0.36 (+1.40%)
PATTON 49.91±0.13 (+3.91%) 49.98±0.38(+0.56%) 52.01±0.50 (+0.23%) 51.07±0.21(+0.53%)
MixGIA 47.49±0.19 (+9.20%) 48.89±0.25 (+2.80%) 50.60±0.19 (+3.02%) 49.73±0.37 (+3.24%)
UniGLM 51.86±0.31 50.26±0.15 52.13±0.34 51.34±0.22
Ogbn-Products SE 53.85±0.17 (+41.99%) 70.52±0.51 (+9.34%) 69.13±0.26 (+12.98%) 69.64±0.17 (+13.47%)
BERT 67.58±0.28 (+13.14%) 74.77±0.87 (+3.13%) 74.09±0.27 (+5.41%) 74.53±0.26 (+6.02%)
GIANT 72.46±0.33 (+5.52%) 69.77±0.42 (+10.52%) 68.69±1.19 (+13.70%) 71.89±0.30 (+9.92%)
PATTON 76.42±0.23 (+0.05%) 77.22±0.34 (-0.14%) 77.81±0.58 (+0.37%) 78.48±0.15 (+0.69%)
MixGIA 71.04±0.38 (+7.63%) 76.13±0.82 (+1.29%) 75.77±0.40 (+3.08%) 76.24±0.33 (+3.65%)
UniGLM 76.46±0.21 77.11±0.41 78.10±0.22 79.02±1.12
History SE 73.19±0.36 (+13.98%) 77.03±0.70 (+7.11%) 77.63±0.22 (+7.29%) 77.83±0.27 (+6.68%)
BERT 78.79±0.31 (+5.88%) 80.33±0.33 (+2.71%) 80.63±0.33 (+3.30%) 80.69±0.22 (+2.90%)
GIANT 81.37±0.32 (+2.52%) 80.88±0.19 (+2.02%) 82.25±0.19 (+1.26%) 81.70±0.26 (+1.63%)
PATTON 82.88±0.24 (+0.65%) 82.41±0.20 (+0.12%) 83.43±0.18 (-0.17%) 82.94±0.07 (+0.11%)
MixGIA 81.47±0.32 (+2.39%) 81.68±0.24 (+1.02%) 82.55±0.23 (+0.90%) 82.26±0.26 (+0.94%)
UniGLM 83.42±0.18 82.51±0.08 83.29±0.17 83.03±0.23
Ogbn-Arxiv SE 64.30±0.09 (+13.56%) 71.74±0.29 (+2.48%) 71.49±0.27 (+4.15%) 74.02±0.18 (+0.19%)
BERT 66.29±0.20 (+10.15%) 72.63±0.31 (+1.23%) 73.33±0.33 (+1.54%) 72.88±0.39 (+1.76%)
GIANT 73.08±0.06 (-0.08%) 73.29±0.10 (-0.31%) 74.59±0.28 (-0.17%) 75.96±0.09 (-2.37%)
PATTON 73.47±0.11(-0.61%) 73.59±0.20(-0.08%) 75.00±0.16(-0.72%) 74.08±0.12(+0.11%)
MixGIA 67.62±0.24 (+7.99%) 73.07±0.30 (+0.62%) 73.70±0.10 (+1.03%) 73.57±0.39 (+0.80%)
UniGLM 73.02±0.11 73.52±0.23 74.46±0.12 74.16±0.51
Table 8: Semi-supervised accuracy results on MLP and state-of-art GNNs with various embeddings for Fitness, Photo, and Ogbn-Products datasets.
Data # Nodes # Edges # Ave Token #classes
PubMed $19,717$ $44,338$ $391$ $3$
Ogbn-Arxiv $169,343$ $1,166,243$ $235$ $40$
Ogbn-Products $54,025$ $74,420$ $163.18$ $47$
Electronics-Computers $87,229$ $808,310$ $117$ $10$
Books-History $41,551$ $400,125$ $302$ $12$
Books-Children $76,875$ $1,631,453$ $280$ $27$
Sports-Fitness $173,055$ $1,946,555$ $30$ $13$
Electronics-Photography $48,362$ $549,290$ $191$ $12$
Table 9: Dataset statistics of eight text-attributed graphs (TAGs).

### v2

@@ 5. Experiments
We aim to answer the following research questions. RQ1: How does UniGLM perform against leading graph embedding models in terms of node classification and link prediction tasks? RQ2: How well does UniGLM transfer in cross-domain and in-domain scenarios? RQ3: How does each component of UniGLM, i.e., sampling strategy, learnable generation and efficient embedding table, contribute to the performance? RQ4: What is the impact of different pre-trained language model backbones on UniGLM?
5.1. Experiments Setup
@@
We train UniGLM on seven TAG datasets (details in Appendix) from two major domains: academic networks and Amazon products. UniGLM is compared with multiple embedding models for node classification and link prediction. Node classification is conducted with three different types of downstream backbones: MLP, GNNs (GCN (Kipf and Welling, 2016a), GraphSAGE (Hamilton et al., 2017), and RevGAT (Li et al., 2021)), and self-supervised graph methods (GraphCL (You et al., 2020) and S2GAE (Tan et al., 2023)), while link prediction uses MLP. Our experiments focus on semi-supervised and transfer learning settings. All baselines are previously discussed in Section 2. Additionally, MixGIANT is implemented by us as a variant of GIANT (Chien et al., 2022) that can be trained on multiple datasets. More details can be found in the Appendix.
5.2. Performance Comparison
Node Classification. To address RQ1, we perform extensive experiments on seven benchmark TAG datasets for node classification. Specifically, we first pre-train UniGLM on these datasets and then use the resultant model to generate node embeddings, which we then use as input features for downstream node classification tasks. The results are shown in Table 2 and Table 3. Comprehensive results are in Appendix. We made the following observations:
① UniGLM significantly outperforms existing graph embedding models in node classification task. Table 2 demonstrates that UniGLM consistently achieves superior performance across most datasets and backbone models, ranking first in 27 out of 28 cases. Although GIANT and PATTON improve over the SE method by leveraging language models and graph structures, UniGLM excels them in most cases. UniGLM’s improvement over GIANT and PATTON indicates that leveraging structural information between graphs is also crucial, further emphasizing the importance of capturing and utilizing these relationships for enhanced performance. ② UniGLM can enhance the performance across different backbones, demonstrating its wide applications. By integrating both structural information within and between graphs, UniGLM can provide a stronger embedding for downstream tasks on a wide range of models.
Dataset Emb Types Photo History Products
GraphCL SE 63.34±0.27 76.14±0.11 64.12±0.05
BERT 55.18±0.89 79.55±0.26 67.52±0.31
GIA 58.39±0.32 68.24±0.09 73.21±0.15
PATTON 75.98±0.14 83.22±0.07 75.91±0.05
UniGLM 78.55±0.10 83.22±0.06 77.08±0.06
Table 3. Semi-supervised accuracy results on classic graph contrastive learning method GraphCL with various embeddings for Photo, History, and Ogbn-Products datasets.
Link Prediction. To address RQ1 for the link prediction task, we train UniGLM on seven datasets with edges in test set masked, using UniGLM as the encoder for the textual attributes of the node, and employ an MLP for downstream evaluation. Detailed split can be found in appendix. The results are illustrated in Figure 2 and more resuls can be found in Appendix. We observe that ③UniGLM exhibits a strong ability in link prediction, surpassing other baselines in most cases. This means that the contrastive learning approach effectively captures and transfers relational patterns between nodes, thereby providing a robust model for various graphs.
Figure 2. Link prediction results in AUC metric.
5.3. Transfer Ability
To answer RQ2, we conduct experiments from two perspectives: in-domain transfer and cross-domain transfer. To assess in-domain transferability, we use the VideoGames co-purchase network as a test case. Detailed settings can be found in the Appendix. Table 4 suggest that ④ UniGLM demonstrate strong in-domain transfer ability in both tasks. The results indicate that incorporating a diverse range of datasets can further enhance the model’s generalization capability.
Link Prediction
SE BERT MixGIANT PATTON UniGLM
AP 68.10±0.12 69.01±0.09 76.82±0.15 75.32±0.12 81.46±0.08
AUC 71.13±0.05 67.51±0.12 75.10±0.09 74.87±0.04 79.58±0.11
Hits@100 0.72±0.04 2.01±0.02 4.04±0.09 4.12±0.12 9.03±0.08
Node Classification
SE BERT MixGIANT PATTON UniGLM
MLP 40.55±0.68 41.88±0.45 45.59±0.36 47.39±0.64 48.70±0.67
GCN 49.28±0.72 49.09±0.54 49.76±0.37 50.16±0.51 51.91±0.63
SAGE 47.20±0.46 48.78±0.56 49.22±0.43 50.32±0.26 52.04±0.43
@@ 5.3. Transfer Ability
Table 4. In-domain transfer ability comparison of UniGLM and baselines on unseen VideoGames dataset for link prediction and node classification tasks.
To assess the cross-domain transfer ability, we use classic ogbn-arxiv as a test case. Detailed settings are reported in the Appendix and Table 5 lists partial results, from which we observe that ⑤UniGLM presents great cross-domain transfer ability, surpassing all baselines in both semi-supervised and few-shot setting. This findings further suggest that language model encoder can learn the structural pattern across a variety of datasets, achieving universal graph embedding.
Dataset Method Accuracy
Semi-Supervised Few-Shot
Ogbn-Arxiv SE 74.02±0.18 15.59±0.83
BERT 66.29±0.20 14.72±1.13
MixGIANT 72.44 ± 0.20 20.16±0.87
PATTON 72.59 ± 0.12 21.08±1.19
UniGraph 72.91±0.42 31.35±1.01
UniGLM 74.60 ± 0.23 33.88±0.87
Table 5. Cross-Domain evaluation of Ogbn-Arxiv under semi-supervised and few-Shot settings.
5.4. Ablation Study
To answer RQ3, we evaluate the performance of the UniGLM after removing or replacing some key components.
The impact of sampling strategy. We designed two variants: one without graph information named “GraphInfo”, where the average degree of the graph is not considered during sampling, and the other is without node information named “NodeInfo”, where the degree of individual nodes is not considered. From Figure 3, we observe that ⑥ Both node degree and graph density information can be used to boost the performance of UniGLM. The possible reason is that domain knowledge is useful for our model to align structure information from various domains in a personalized fashion.
Figure 3. Ablation Study: the impact of different sample strategies. Results are the average of MLP, GCN and SAGE backbones.
The imapct of learnable positive generation module. We introduce two variants, “w/o W”, which fixes the weight instead of make it learnable. The second approach is “a_sim”, which uses the weights to aggregate the similarities between the positive samples and the center node, rather than generating the embeddings of the positive samples themselves before computing similarity. The results are shown in Figure 4, from which we observe that ⑦ Our learnable positive generation module performs consistently better than other variants. We attribute this improvement to the selection mechanism inherent in the positive embedding generation process, which has the potential to constitute more informative positive samples. To further verify this, we visualize the positive candidates and the generated positive embedding in Figure 5. We observe that the generated positive samples often lie in the middle of corresponding positive candidates, leading to an informative positive sample for constrastive loss compared with original ones.
Figure 4. The impact of learnable positive generation scheme on UniGLM. The results are averaged values across three backbones: MLP, GCN, and SAGE.
Figure 5. Visualization of the generated positive sample and the corresponding positive candidates on History dataset.
How effective is the proposed embedding table module? We study the influence of the lazy contrastive module in Section 4.3. From the efficiency perspective, as shown in Table 1, ⑧ Our lazy contrastive module can significantly accelerate the training speed, since it avoids the calculations of positive samples on-the-fly. More interestingly, we empirically observe that this lazy updating does not impair the model performance as expected. In contrast, it can also boost our model performance as shown in Figure 6. The feasible reason is that this lazy update rule allows our model to remember the most recent record during training, thus greatly reducing the risk of deviating from the previous step based solely on the current mini-batch.
Figure 6. The impact of lazy contrastive module on UniGLM. Results are the average of MLP, GCN and SAGE backbones.
To answer RQ4, we test the performance of UniGLM under three different PLM backbones: MPNET(Song et al., 2020), BERT and DeBERTa(He et al., 2020b). The results are summarized in Table 6, from which we observe that ⑨ The performance of UniGLM varies across various PLM backbones, yet it can achieve consistently better results using Sentence-BERT.
Model ogbn-arxiv Children Computers
Sentence-mpnet 74.34 ± 0.09 52.36 ± 0.26 76.40 ± 0.11
BERT 73.51 ± 0.11 51.92 ± 0.28 75.43 ± 0.08
DeBERTa 73.23 ± 0.23 51.98 ± 0.42 74.82 ± 0.16
Table 6. The impact of PLM backbones on UniGLM.
6. Conclusion
@@
In this paper, we introduce UniGLM, a unified graph embedding framework aimed at fine-tuning pre-trained language model on a set of TAG datasets from various domains. UniGLM features two main innovations: (1) an adaptive and learnable positive scheme that selects positive samples using node’s local, global, and domain knowledge, which is subsequently used for informative positive sample generation, and (2) a dynamic embedding table that efficiently encodes these samples on-the-fly to speed up training. We validate UniGLM across diverse TAGs from different domains, where it consistently surpasses existing methods in node classification, link prediction, demonstrating its superior transfer ability, effectiveness and efficiency.